The grade point average of the students at BLK is 2.80 with a standard deviation of 0.84. The grade point average of students at SJK is 2.4 with a standard deviation of 0.84. Which university shows a more dispersed grade distribution? Explain you reasoning with statistical evidence.
Answer #1

Solution :

Given that , For BLK , \mu = 2.8 and \sigma = 0.84

and for SJK , \mu = 2.4, \sigma = 0.84

We will compute Coefficient of variation to check the dispersion.

CV for BLK is,

CV = \frac{\sigma }{\mu } = (0.84/2.8) = 30

BLK's coefficient of variation is 30% .

CV for SJ/K is,

CV = \frac{\sigma }{\mu } = (0.84/2.4) = 0.35

SJK's coefficient of variation is 35% .

SJK's coefficient of variation = 35%. BLK's coefficient of variation = 30%. Therefore, SJK has a more dispersed grade distribution.